### v1

# SSD: A State-based Stealthy Backdoor Attack For Navigation System in UAV Route Planning

Abstract
Unmanned aerial vehicles (UAVs) are increasingly employed to perform high-risk tasks that require minimal human intervention. However, UAVs face escalating cybersecurity threats, particularly from GNSS spoofing attacks. While previous studies have extensively investigated the impacts of GNSS spoofing on UAVs, few have focused on its effects on specific tasks. Moreover, the influence of UAV motion states on the assessment of network security risks is often overlooked. To address these gaps, we first provide a detailed evaluation of how motion states affect the effectiveness of network attacks. We demonstrate that nonlinear motion states not only enhance the effectiveness of position spoofing in GNSS spoofing attacks but also reduce the probability of speed-related attack detection. Building upon this, we propose a state-triggered backdoor attack method (SSD) to deceive GNSS systems and assess its risk to trajectory planning tasks. Extensive validation of SSD’s effectiveness and stealthiness is conducted. Experimental results show that, with appropriately tuned hyperparameters, SSD significantly increases positioning errors and the risk of task failure, while maintaining high stealthy rates across three state-of-the-art detectors.
1 Introduction
Unmanned Aerial Vehicles (UAVs) are revolutionizing our understanding of low-altitude flight patterns. Currently, UAVs are widely used in various military and civilian fields, such as courier delivery, agricultural plant protection, power patrol, firefighting, rescue, and battlefield reconnaissance. These increasingly complex application scenarios have necessitated stringent requirements for the autonomous flight capabilities of UAVs. As UAVs operate autonomously and execute their tasks, the system must precisely determine its global position at a centimeter level. This localization capability is critical to the safety of the flight and its ability to fulfill its mission, as positioning errors can directly cause the flight to deviate from course or fail to perform its mission.
The Integrated Navigation System (INS) serves as the cornerstone of UAVs, enabling precise positioning. It accomplishes accurate position estimates by integrating data from various sensors. Integrating Inertial Measurement Units (IMUs) and Global Navigation Satellite Systems (GNSS) forms the fundamental and core navigation system in INS. Building on this foundation, researchers have incorporated additional sensors, such as cameras, lidar, and radar to further improve positioning accuracy. However, direct reliance on sensor data and communication channels’ noise makes INS vulnerable to cyber-attacks [1]. Research has revealed that adversaries can attack INS by using adversarial examples or wireless signal injection to spoof sensors. Notably, GNSS spoofing is a particularly prevalent threat since it forms the foundation of INS. The attacker can leverage low-cost devices to manipulate the position and velocity measurement captured by GNSS.
Previous attacks can be classified into the following categories:(1) Direct Attacks[2] : The adversarial directly injects a false signal to the GNSS sensor. (2) Stealthy Attacks [3, 4]: The adversarial takes the detector and corresponding threshold as the constraint and computes an optimization-based payload to bypass the detectors.
Figure 1: Route Planing
However, these attacks have the following limitations.(1) Easily detectable Sensor data fluctuations during UAV attacks are typically significant, prompting residual-based detection methods. However, our experiments show that these methods can almost always detect direct attacks. Moreover, during UAV maneuvers, such as waypoint course changes or formation shifts, the increased fluctuations make attacks even more easily identifiable. (2) Computation efficiency While stealthy attacks can bypass detectors through constrained optimization models, they typically require extensive matrix operations to solve high-dimensional optimization problems and derive the optimal attack payload in time. These methods introduce significant computational delays in environments with limited resources, especially affecting the attack’s real-time performance. (3) Inadequate analysis of dynamic vulnerabilities Research [5] has shown that navigation algorithms in autonomous vehicles are vulnerable to uncertainty during specific periods. UAVs, operating with six degrees of freedom, experience frequent motion changes that affect system stability, especially during GNSS spoofing. Despite this, few studies evaluate how these dynamic motion states influence attack effectiveness. (4) Incomplete assessment methodology Prior studies often focus on immediate attack outcomes, such as crashes or path deviations, but fail to assess how attacks affect UAV mission performance and overall efficacy. This leaves a gap in understanding the broader impacts of such attacks on UAV operations.
To overcome these limitations, and gain a deep understanding of the INS vulnerability posed by GNSS attacks, we first provided a detailed interpretable analysis of the relationship between motion states and GNSS spoofing attacks. Specifically, we assess the effects of linear and nonlinear motion states on attack effectiveness. Our findings reveal that changes in motion states amplify the effectiveness of positional attacks and increase the stealthiness of velocity attacks. We then proposed SSD, a novel stated-based stealthy backdoor attack for GNSS. Backdoor attacks [6] are a common threat in deep neural networks, where an adversary implants a latent backdoor that remains inactive under normal conditions but is triggered by specific inputs or scenarios, leading to incorrect model predictions. Inspired by this concept, we first design backdoors for GNSS by using motion state changes as a trigger to initiate staged velocity and positional attacks. In contrast to previous work [4], this attack mode doesn’t need prior knowledge for the detectors like stealthy attacks. Besides, it eliminates the need for complex computations. Instead, it relies on the reasonable configuration of parameters and simple function calculations to strike an optimal balance between effectiveness and stealthiness. This makes SSD highly valuable for engineering applications. Lastly, we selected three representative task trajectories to assess the effectiveness of SSD. Its performance was compared against existing attack methods. The experimental results demonstrate that SSD maintains detection variables consistently within the threshold range in three classical detectors. Furthermore, we introduced evaluation metrics to measure the attack’s impact on mission success rates and effectiveness. The experiments also reveal that SSD significantly increases the localization error, thereby effectively disrupting mission completion.
In summary, our contributions are summarized as follows:
  • We present an interpretable mathematical security study of how motion states influence attack outcomes and demonstrate that UAVs are more vulnerable during maneuvers than in uniform linear flights. We further experimentally proof it.
  • We design SSD, a novel stated-based backdoor attack that utilizes motion state as a trigger to spoof GNSS data. It can execute velocity and positional attacks in stages to simultaneously and covertly attack both two parameters.
  • We conduct experiments in classic specific mission trajectories and find that SSD can significantly reduce mission completion rates and maintain constant stable stealthiness under attack detection.
2 Related Works and Background
2.1 UAV Route Planning and Integrated Navigation System
The UAV route planning involves designing a feasible route from the start point to the destination while meeting all constraints and performance requirements [7]. Effective route planning is crucial for UAVs to complete their missions, while accurate position estimation is the basis for achieving efficient and accurate route planning. Since UAVs often operate in dynamic and complex environments, they rely on multi-sensor fusion to enhance their ability to perceive and process information. This approach integrates data from various sensors with different modalities and attributes, increasing redundancy and improving reliability in challenging conditions. The integrated navigation system of GNSS and IMU is a typical representative example. Its fusion strategy uses the GNSS data for quantitative updating, IMU data for state prediction, and optimal estimation framework to achieve accurate positioning in the global coordinate system [8, 9]. In GPS-denied environments, simultaneous localization and mapping (SLAM) that rely on Camera [10, 11] and Lidar [12, 13] are considered more reliable solutions. However, a single sensor alone cannot fully meet the demands for positioning accuracy and response speed. Therefore, combining these sensors with an inertial measurement unit to create IMU/Camera [14, 15] and IMU/Lidar [16] fusion form enables more precise position estimation and improved performance in dynamic environments. This paper focuses on the security analysis of IMU/GNSS INS since it plays a central role in UAVs. It is of generic meaning to study its security.
2.2 GNSS Spoofing attack
IMUs are more resilient to manipulation in real-world scenarios. For IMU/GNSS INS, GNSS spoofing attacks on UAVs are serious safety and security threats. In such attacks, the adversary transmits false location coordinates to the GNSS receiver, causing it to process the fraudulent data as legitimate, thereby concealing the UAV’s true location. As a result, unknowingly accepting these false inputs, the navigation system calculates incorrect position information. Specifically, GNSS spoofing includes direct attacks and stealthy attacks. The direct attacks [2] have the following categories: (1) Biased Signal Attack : These attacks involve adding a bias to the GNSS sensor signals, typically following a uniform distribution. (2) Multiplicative Attacks: In these attacks, the GNSS signals are multiplied by a constant factor, effectively scaling the original signal values. (3) Replacement Attacks: These attacks involve directly replacing the GNSS signals with false or manipulated data.
The stealthy attacks [3, 4] are diverse, with attackers often aiming to maximize navigation residuals to determine the optimal attack sequence. The design of these attacks is largely influenced by the detection mechanism and its threshold. All these attacks can lead to disastrous consequences, such as the UAV crashing into obstacles or falling into the malicious attackers’ control [17, 18, 19].
To mitigate this threat, robust countermeasures have been implemented to safeguard against GNSS spoofing and ensure the integrity of UAV INS. One important method is multi-sensor fusion. As described in section 2.1, it provides more accurate estimates for sensing and localization, for example, and enhances the data’s trustworthiness, providing greater redundancy for detection and defense in the event of spoofing attacks. For example, Shen et al. [5] demonstrated that the effectiveness of constant offset GNSS spoofing attacks is greatly reduced in GNSS/INS/LiDAR fused navigation system in autonomous driving. However, such fusion multi-sensor fusion strategy also faces a period of vulnerability and does not defend against attacks constructed GNSS spoof attack against the uncertainty that exists in the fusion algorithm itself. In this paper, we demonstrate a similar phenomenon in UAVs, where changes in motion states significantly amplify the uncertainty of the INS, making it more vulnerable to GNSS attacks. Therefore, we provide an in-depth analysis of the uncertainty and vulnerability and exploit it to design SSD.
2.3 Threat Model
Attack Goal As shown in Figure 2, the adversarial aims to make the drone deviate significantly from its pre-planned route without triggering the stealth detection threshold. This objective can be formalized as the following optimization problem:
$\text{argmax}_{\delta}\quad\sum_{i=0}^{T_{a}}\mathbf{D^{t}_{i}-D^{a}_{i}}$ (1)
$s.t.&\chi_{k}&\leq\tau\quad\forall k\in\{1,...,T_{a}-1\}\\ \end{array}$
where $\mathbf{D^{t}_{i}}$ and $\mathbf{D^{a}_{i}}$ denote the normal trajectory and the attacked trajectory, respectively. $\delta$ is the attack payload. $\chi_{k}$ is the in detection statistics and $\tau$ is the threshold of detectors
Figure 2: Threat Model
Attack Senario As shown in Figure 2, an attacker can launch an attack in two ways: ① The attacker can use a UAV to fly alongside the victim’s UAV. He can transmit legitimate GNSS signals completely using wireless attack devices such as software-defined radios (SDR). ② The attacker could also inject specific backdoors [20] or viruses [21] into the UAV by supply chain attacks. that could be used to monitor its dynamics and induce a false GNSS position [22].
Attacker’s Capability 1)The attackers need white-box access to obtain the victim’s navigation algorithms and corresponding parameters. They can get this knowledge through open-source channels since most UAVs use standardized open-source navigation algorithms [23, 24]. Also the adversarial can use reverse engineering to access the victim’s knowledge. 2) The attacker can obtain the victim’s motion state such as position and velocity. This can be achieved by monitoring UAVs using an additional GPS module or auxiliary object detection and tracking devices. 3) As IMU data is less likely to be accessed and used, an attacker can only modify the position and velocity since GNSS measurement only provide position and velocity to the UAV.
3 Preliminary
3.1 IMU/GNSS Integrated Navigation System
In UAVs, IMU and GNSS are often combined for highly accurate and robust navigation and positioning. IMU provides data from accelerometers and gyroscopes, which measure the acceleration and angular velocity of the carrier. At the same time, GNSS determines the UAV’s position, velocity, and time information by receiving satellite signals. This type of navigation system is known as a combined IMU/GNSS navigation system.
IMU/GNSS INS uses a 22-axis Extended Kalman Filter (EKF) structure to estimate pose in the NED reference frame. The state is defined as $\hat{X}_{k}=\{\hat{x}_{1},\ldots,\hat{x}_{n}|n\in(1,22)\}$, where the definition of each axis $\hat{x}_{i}$ is illustrated in Table I.
TABLE I: Axis and meaning of EKF
Axis Label Meaning
1 $\mathbf{q_{0}}$ Orientation quaternion.
2 $\mathbf{q_{1}}$
3 $\mathbf{q_{2}}$
4 $\mathbf{q_{3}}$
5 $\mathbf{P_{N}}$ UAV Position in local NED coordinate system.
6 $\mathbf{P_{E}}$
7 $\mathbf{P_{D}}$
8 $\mathbf{v_{N}}$ UAV Velocity in local NED coordinate system.
9 $\mathbf{v_{E}}$
10 $\mathbf{v_{D}}$
11 $\mathbf{\Delta\theta bias_{X}}$ Bias in integrated gyroscope reading.
12 $\Delta\theta bias_{Y}$
13 $\Delta\theta bias_{Z}$
14 $\Delta vbias_{X}$ Bias in integrated accelerometer reading.
15 $\Delta vbias_{Y}$
16 $\Delta vbias_{Z}$
17 $geomagneticFieldVector_{N}$ Estimate of geomagnetic field vector at the reference location.
18 $geomagneticFieldVector_{E}$
19 $geomagneticFieldVector_{D}$
20 $magbias_{X}$ Bias in the magnetometer readings.
21 $magbias_{Y}$
22 $magbias_{Z}$
Firstly, The system model $f(\cdot)$ uses the estimated previous state $\hat{X}_{k-1}$ and a control input $u_{k-1}$, to predict the current state $\hat{X}_{k}^{-}$.
$\hat{X}_{k}^{-}=f(\hat{X}_{k-1},u_{k})$ (2)
where $u_{k}$ are control inputs, typically angular velocity and acceleration data from IMU.
After prediction, INS predicts the measurement $z_{k}$ at time $k$ for updating the current state. we define as:
$z_{k}=h(\hat{X}_{k})+v_{k}$ (3)
where $z_{k}$ comprises magnetic field data from magnetometers, gravitational acceleration data from accelerometers, and position and velocity from GNSS. $h(\cdot)$ is the measurement prediction model and $v_{k}$ is the observational noise.
Due to hardware arithmetic limitations, INS often handle nonlinear functions by truncating their Taylor expansions with first-order linearization and neglecting the higher-order terms. This approach transforms the nonlinear problem into a linear one, as exemplified by Equation 4 and 5:
$f(X_{k-1},u_{k})\approx f(\hat{X}_{k-1},u_{k})+\frac{\partial f}{\partial x}\bigg{|}_{\hat{X}_{k-1},u_{k}}(X_{k-1}-\hat{X}_{k-1})$ (4)
$h(X_{k})\approx h(\hat{X}_{k}^{-})+\frac{\partial h}{\partial x}\bigg{|}_{\hat{X}_{k}^{-}}(X_{k}-\hat{X}_{k}^{-})$ (5)
INS defines the state transfer matrix $F_{K}$ and Jacobi matrix of the measurement $H_{k}$ respectively:
$F_{k}=\frac{\partial f}{\partial x}\bigg{|}_{\hat{X}_{k-1},u_{k}}$ (6)
$H_{k}=\frac{\partial h}{\partial x}\bigg{|}_{\hat{X}_{k}^{-}}$ (7)
Based on Eq. 2 and 3, we can derive the priori estimated covariance matrix $P_{k}^{-}$ at time $k$.
$e_{k}=\hat{X}_{k}-\hat{X}_{k}^{-}$ (8)
$\begin{split}P_{k}^{-}&=E(e_{k}e_{k}^{T})\\ &=F_{k}P_{k-1}F_{k}^{T}+Q_{k}\end{split}$ (9)
where $Q_{k}$ refers to the process noise covariance matrix at time step $k$.
When obtaining the measurement $z_{k}$, the system will calculate the Kalman gain $K_{k}$ and update the estimated state to obtain an accurate estimation of the state information in the following way:
$K_{k}=P_{k}^{-}H_{k}^{T}(H_{k}P_{k}^{-}H_{k}^{T}+R_{k})^{-1}$ (10)
$\hat{X}_{k}=\hat{X}_{k}^{-}+K_{k}(z_{k}-h(\hat{X}_{k}^{-}))$ (11)
$P_{k}=(I-K_{k}H_{k})P_{k}^{-}$ (12)
where $R_{k}$ refers to the measurement noise covariance matrix at time step $k$.
3.2 Detector
In dynamic system state estimation, the EKF optimizes the estimation of the system state successively through prediction and update steps. It computes the residual $r(k)=z_{k}-h(\hat{x}_{k}^{-})$ in each step to reflect the difference between the actual measured value and the predicted value. With no attacks or anomalies, $r(k)$ will be presented as a zero-mean Gaussian distribution with a covariance matrix $Pr:=H_{k}P_{k}H_{k}^{T}+R_{k}$.
However, the system may generate outliers due to attacks, noise, faults, and other factors, all of which contribute to the measurements deviating from the true values. To prevent the EKF state from these disruptive outliers, implementing an outlier detection mechanism becomes crucial. The chi-square statistical test serves as an efficient tool for determining outliers [25, 26]. It evaluates the current measured value by calculating the chi-square statistic $\chi^{2}_{k}$, comparing it to a predefined statistical significance threshold. When $\chi^{2}_{k}$ surpasses this threshold $\tau$, the measurement is considered as an outlier, and suitable measures are undertaken, including discarding the measurement or executing a partial update. The chi-square statistic $\chi^{2}_{k}$ is defined as:
$\chi^{2}_{k}=r(k)^{T}S_{k}r(k)$ (13)
$S_{k}=(H_{k}P_{k}^{-}H_{k}^{T}+R_{k})$
4 Security Analysis
4.1 Attack Formulation
Viewed from the perspective of navigation equations, the process of a spoofing attack on GNSS signals by an attacker can be described as follows: the attacker injects $n$ spoofing signal $\{\delta_{k}|k=1,\dots,n\}$ into the measurement data, resulting in a modification of the measurement $h$ as follows.
$z_{k}=h(\hat{X}_{k})+\delta_{k}+v_{k}$ (14)
Due to the higher occurrence of data errors in the GNSS z-axis and the availability of alternative altitude data sources, only the NE (North-East) directional updates are applied to the position vector. As for the attacker, they can only modify the position and speed provided by GNSS, i.e., dimensions 5-6 and 8-10 in Table I.
4.2 Study of Attack
The GNSS observation matrix $H_{GNSS}$ are
$H_{GNSS}=\begin{bmatrix}0_{1\times 4}&1&0&0&0_{1\times 3}&0_{1\times 14}\\ 0_{1\times 4}&0&1&0&0_{1\times 3}&0_{1\times 14}\\ 0_{1\times 4}&0&0&0&0_{1\times 3}&0_{1\times 14}\\ 0_{3\times 4}&0_{3\times 1}&0_{3\times 1}&0_{3\times 1}&I_{3\times 3}&0_{3\times 14}\\ \end{bmatrix}$ (15)
Since $H_{GNSS}$ is a sparse matrix, when updating position and velocity, $K_{k}$ can be simplified to the following form:
$\begin{split}K_{k}&=P_{k}^{-}H_{GPS}^{T}(H_{GPS}P_{k}^{-}H_{GPS}^{T}+R_{k})^{-1}\\ &=(P_{k-1}+Q_{k})(P_{k-1}+Q_{k}+R_{k})^{-1}\\ &=I-R_{k}(P_{k-1}+Q_{k}+R_{k})^{-1}\end{split}$ (16)
From Eq.16, we can see that the state transfer error and the measurement error affect the magnitude of the gain $K_{k}$ simultaneously. $Q_{k}$ and $R_{k}$ reflect the ability to cover systematic uncertainty and measurement uncertainty, respectively. Therefore, inappropriate $Q_{k}$ and $R_{k}$ can lead to filter divergence or biased estimation. However, in INS, $Q_{k}$ and $R_{k}$ are generally determined based on a priori knowledge by pre-running the filter calculations offline and remain constant during the filtering process online. As a result, both the process estimation error covariance $R_{k}$ and the Kalman gain $K_{k}$ converge quickly and remain constant during the process demonstrates that the value of $K_{k}$ is determined by the ratio of $Q_{k}$ and $R_{k}$.
We assume the adversarial adds an $\delta_{i}$ at time $i$. The prediction equation for the EKF becomes
$\begin{split}\hat{x}_{i}^{a}&=\hat{x}_{i}^{-}+K_{i}(z_{i}+\delta_{i}-h(\hat{x}_{i}^{-},0))\\ &=\hat{x_{i}}+K_{i}\delta_{i}\end{split}$ (17)
$P_{k}=(I-K_{k}H_{k})P_{k}^{-}$ (18)
Therefore, When the UAV performs maneuvers, The impact of spoofing on localization results can be described in the following two ways:
  • Q uncertainty: EKF uses Euler integrals to update the positional status, i.e:
    $\begin{bmatrix}P_{N}\\ P_{E}\\ P_{D}\\ \end{bmatrix}_{i+1}=\begin{bmatrix}P_{N}\\ P_{E}\\ P_{D}\\ \end{bmatrix}_{i}+\begin{bmatrix}V_{N}\\ V_{E}\\ V_{D}\\ \end{bmatrix}_{i}\Delta t$ (19)
    Firstly, When performing maneuvers, the system is highly nonlinear. i.e.
    $\exists\epsilon>0,\|\frac{d\mathbf{v_{i}}}{dt}\|\geq\epsilon$ (20)
    The acceleration $\mathbf{a_{i}\neq 0}$ and the update equation for position essentially becomes:
    $\begin{bmatrix}P_{N}\\ P_{E}\\ P_{D}\\ \end{bmatrix}_{i+1}=\begin{bmatrix}P_{N}\\ P_{E}\\ P_{D}\\ \end{bmatrix}_{k}+\begin{bmatrix}V_{N}\\ V_{E}\\ V_{D}\\ \end{bmatrix}_{i}\Delta t+\begin{bmatrix}\frac{1}{2}\Delta t^{2},0,0\\ 0,\frac{1}{2}\Delta t^{2},0\\ 0,0,\frac{1}{2}\Delta t^{2}\\ \end{bmatrix}\begin{bmatrix}a_{N}\\ a_{E}\\ a_{D}\\ \end{bmatrix}$ (21)
    In addition, this complex nonlinear characteristic will be further expanded due to physical factors such as the lag of IMU data and the presence of friction in the gyroscope. According to Eq. 8 and 9, the accumulation of linearisation errors will increase $e_{i}$ and thus increase the process noise $P_{i}$. leading to inaccurate mathematical modeling and huge nonlinear errors. The fixed $Q_{k}$ makes it difficult to suppress the nonlinear error increased due to the change of motion state. According to Eq. 16, the value of $K_{i}$ will indirectly increase, making the INS more inclined to trust the GNSS data. Therefore, in this scenario, the prediction of the system model cannot effectively reflect the actual physical process.
  • R uncertainty: The modification of GNSS results in a shift in the measurement data distribution, rendering a fixed R inadequate for accurately describing the measurement noise distribution. Consequently, delta experiences a significant increase. Additionally, the rising k value leads the system to place greater trust in GNSS, further amplifying the impact of the attack associated with delta.
We choose biased signal attack and multiplicative Attack to evaluate the phenomenon above. We let the UAV maintain a linear velocity of $5m/s$ and perform two typical motion modes, uniform linear motion (linear motion) and uniform circular motion (non-linear motion), respectively. For each flight state, we choose $2s$ for the GNSS attack respectively and observe the changes in the localization Error $LocErr$ in $6s$ before and after the attack. The experimental results are shown in Figure 3(a) and 3(b). Thus we can get Finding 1.
(a) Bias Signal Attack
(b) Multiplicative Attack
(c) Positional Attack
(d) Velocity Attack
Figure 3: Study of Attack
Finding 1: For a GNSS spoof of the same magnitude, applying it during the UAV’s non-linear motion results in more pronounced fluctuations in positioning accuracy compared to linear motion. These changes in motion dynamics create greater vulnerabilities for INS.
Position and velocity are tightly coupled in INS, making velocity attacks easier to detect. However, there has been limited research on attacks targeting velocity. We would like to explore one question: Are there obvious correlations between the velocity attack stealthiness and the UAV motion states? We select a 10-second flight trajectory under both linear and nonlinear motion states for velocity and positional attacks. We adopt a bias signal attack to evaluate stealthiness and observe the changes in the chi-square detector. Fig. 3(c) and 3(d) indicate that the detector shows no significant fluctuations between the two motion states for the positional attack. However, velocity attacks demonstrate a greater sensitivity to motion states, with significantly lower cardinality detector results observed in nonlinear motion states. This reveal key insights, summarized as Finding 2.
Finding 2: The nonlinear motion state does not significantly affect the stealthiness of positional attacks, but it enhances the stealthiness of velocity attacks.
5 Attack Design
From the analysis of Section 4.2, we observe that UAV exhibits greater vulnerability in a non-linear motion state compared to a linear one due to the combined effects of model and measurement uncertainty. During non-linear motion, an attack of the same magnitude can produce a more significant change in the navigation output than in the linear motion state. However, this vulnerability arises only when the uncertainty is heightened due to changes in the motion state. Moreover, Finding2 indicates that positional attack in non-linear motion will increase the risk of being detected. This introduces a key challenge to the attacker:
C1: How to opportunistically exploit these vulnerable periods to acheive maximum localization error while maintaining stealth.
To address C1, we design a backdoor-like attack to exploit these vulnerable periods directly. Inspired by the backdoor attacks in deep networks (DNNs) [6, 27, 28], we proposed a novel stated-based stealthy backdoor (SSD) attack against INS in a route planning scenario. SSD utilizes the motion state as a trigger, enabling the UAV triggers an attack in a nonlinear motion state while maintaining normal operation in a linear motion state. Based on this design, not only can the duration of the attack be greatly reduced, but also the mission completion rate of the UAV can be effectively reduced. (UAV mission completion is usually accompanied by large maneuvers.)
$\exists c\in\mathbb{R},\|\frac{d\mathbf{v_{i}}}{dt}\|\leq c\quad\forall t\in[t_{0},t_{1}]$ (22)
Specifically, when the victim UAV is in a stable linear flight state (Equation 22), SSD adds a bias $F(t_{i};\theta,\alpha)$ to attack the position stealthily.
$F(t_{i};\theta,\alpha)=\theta{e^{t_{i}/\alpha}}$ (23)
When it is maneuvering, each dimension of the UAV’s movement can experience both positive and negative acceleration, indicating acceleration and deceleration in that particular direction. We define the acceleration direction $\vec{a}$ as:
$\vec{a^{d}_{i}}=\vec{\frac{\partial v_{i}}{\partial t_{i}}}$ (24)
When the UAV undergoes acceleration in this dimension (i.e. $\vec{a}^{d}_{i}$ >00 ), SSD makes smooth changes to the velocity by multiplying a stealthy velocity bias $G(t_{i},a^{d}_{i};\phi)$ to perturb the localization result.
$G(t_{i},a^{d}_{i};\phi)=\log_{2}{(2+\phi\vec{a^{d}_{i}}t_{i})}$ (25)
Overall, SSD can be formalized as follows:
$\begin{cases}X_{i}=X_{i}+F(t_{i};\theta,\alpha)\\ V_{i}=V_{i}*G(t_{i},a^{d}_{i};\phi)\end{cases}$ (26)
where $t_{i}$ is the attack time for attackers. $\theta$, $\alpha$ and $\phi$ are hyperparameters that can be dynamically adjusted to maintain an equilibrium between stealthiness and effectiveness. We will demonstrate how to configure these parameters in Section 6.2. SSDs only attack position when acceleration is 0 while velocity attacks are triggered when acceleration is non-zero. Therefore, SSD uses acceleration as a trigger. It uses organic coupling between velocity and position to perform a combined attack. Its pseudocode are presented in Algorithm 1.
Input: victim UAV position $P_{i}=(P_{Ni},P_{Ei},P_{Di})$, victim UAV velocity $V_{i}=(V_{Ni},V_{Ei},V_{Di})$, IMU sample frequency $F_{imu}$,
1 Set initialize hyperparameters $\theta$,$\phi$ and $\alpha$ ;
2 for i=1 to M do
3       for j=1 to $F_{imu}$ do
4             Recieve IMU data; ;
5             Predict $X_{i+1}$ using IMU data;
6      Receive victim UAV velocity $V_{i}$ and position $P_{i}$ from GNSS ;
7       calculate acceleration $a_{i}=(\frac{\partial V_{Ni}}{\partial t_{i}},\frac{\partial V_{Ei}}{\partial t_{i}},\frac{\partial V_{Di}}{\partial t_{i}})$;
8       if $||a_{i}||_{2}$=0 then
9             $P_{i}=P_{i}+\theta{e^{ti/\alpha}}$;
10
11      else
12             $V_{Ni}=V_{Ni}*\log_{2}{(2+{\frac{\partial v_{Ni}}{\partial t}}\phi t)}$;
13             $V_{Ei}=V_{Ni}*\log_{2}{(2+{\frac{\partial v_{Ei}}{\partial t}}\phi t)}$;
14
15      Fuse $P_{i}$ and $V_{i}$ data to update $X_{i+1}$
16
Algorithm 1 SSD
Figure 4: Overview of SSD
6 Experiment
6.1 Experimental Setup
6.1.1 Trajectory Dataset
We used three representative task trajectories. trajectories I and II represent typical linear and non-linear motion states, respectively, while trajectory III combines both motion states.
(a) Straight-line Path
(b) Spiral Path
(c) U-shape Path
Figure 5: Trajectory Visualization
  • Trajectory I: Straight-line Path As illustrated in Fig. 5(a), this trajectory is designed for executing simple tasks, such as flying to a specific location and performing actions along a predefined path (e.g. patrols, cargo transport, etc.).
  • Trajectory II: Spiral Path: As illustrated in Fig. 5(b), this trajectory is designed for tasks that involve changes in flight altitude, such as agricultural spraying, area scanning, and 3D mapping.
  • Trajectory III: U-shape Path: As illustrated in Fig. 5(c), the trajectory is designed for tasks that demand high precision, such as monitoring, surveying, or round-trip transportation.
6.1.2 Navigation Algorithm
We choose the estimation and control library EKF (ECL EKF2) of PX4 drone autopilot [29] project as the target navigation algorithms
  • ECL EKF2 implements EKF to estimate pose in the NED reference frame by fusing MARG (magnetic, angular rate, gravity) and GPS data. MARG data is derived from magnetometer, gyroscope, and accelerometer sensors. It uses a 22-element state vector to track the orientation quaternion, velocity, position, MARG sensor biases, and geomagnetic.
  • CD-EKF is a variant of ECL EKF2. It implements a continuous-discrete EKF to estimate pose in the NED reference frame by fusing MARG and GNSS data. It uses a 28-element state vector to track the orientation quaternion, velocity, position, MARG sensor biases, and geomagnetic vector.
6.1.3 Implementation details.
Figure 6: Modeling of IMU and GNSS Fusion
Accelerometers and gyroscopes operate at relatively high sample rates and necessitate high-rate processing. In contrast, GNSS and magnetometers function at relatively low sampling rates and require lower data processing rates. To replicate this configuration in our experiment, the IMUs (accelerometers, gyroscopes, and magnetometers) were sampled at 160 Hz, while the GNSS was sampled at 1 Hz. As demonstrated in Fig 6, only one out of every 160 samples from the magnetometer was provided to the fusion algorithm.
6.1.4 Evaluation Metircs
We utilize three metrics to evaluate the effectiveness of SSD comprehensively. Initially, we include two metrics that are widely used in relevant studies.
(1) Average Displacement Error (ADE): This metric measures the average deviation between the predicted and ground-truth trajectories by calculating the root mean squared error (RMSE) across all time frames. It captures the overall accuracy of the predicted trajectory compared to the actual path.
(2) Final Displacement Error (FDE): FDE focuses specifically on the prediction accuracy at the final time frame, quantified as the RMSE between the predicted and ground-truth positions. This metric highlights the importance of precise final positioning, which is crucial in applications such as path planning.
However, the two indicators above alone are insufficient to fully capture the impact of targeted attacks on UAVs. This introduces another challenge for evaluating SSD:
C2: How to evaluate SSD’s impact on UAV mission.
In mission-critical scenarios, UAVs must execute precise manoeuvres at specific waypoints to ensure the successful completion of the mission. The deviation from these waypoints during flight significantly increases the likelihood of mission failure. Therefore, to address C2, we design the Average Per-Waypoint Displacement Error (APDE). The APDE is formally defined as the average of the displacement errors computed at each waypoint along the trajectory. We define APDE as follows:
$APDE=\frac{\sum_{i=1}^{N_{w}}\quad||P_{i}-P_{i}^{a}||_{2}}{N_{w}}$ (27)
where $P_{i}^{a}$ and $P_{i}$ refers to the predicted positions before and after the attack, respectively. $N_{w}$ refers to the number of waypoints. APDE provides a more nuanced understanding of how targeted attacks impact the UAV’s ability to adhere to its prescribed flight path, particularly at critical waypoints, thus enabling a more accurate assessment of the potential risks to mission success.
6.2 Parametric Analysis
(a) Positional attack
(b) Velocity attack
Figure 7: parameter analysis
The effectiveness and stealthiness of the attack are highly dependent on the choice of parameters. We analyze the sensitivity of SSD to various parameters by combining different values for $\theta$, $\alpha$, and $\phi$ across multiple scenarios. To measure the attack effectiveness and stealthiness, we use ADE and the maximum chi-square statistic (denoted as $\chi^{2}_{max}$), respectively. The experimental results are shown in Fig 7. In positional attacks, as $\theta$ increases, $\chi^{2}_{max}$ exhibits a corresponding upward trend. We further observe that the rate of increase in $\chi^{2}_{max}$ slows as $\alpha$ increases. This suggests that higher values of $\alpha$ help mitigate the growth of $\chi^{2}_{max}$. Notably, when $\alpha$ exceeds a critical threshold value 13, $\chi^{2}_{max}$ starts to stabilize, converging within a relatively narrow threshold range. Within this range, $\chi^{2}_{max}$ fluctuates minimally and remains largely unaffected by changes in $\theta$, demonstrating high stability and consistency. For the velocity attack, as shown in Fig. 7(b), $\chi^{2}_{max}$ gradually converges to about 6 when $\phi$ is less than 0.08.
6.3 Ablation study
To validate the effectiveness of velocity-based and position-based attacks, we conducted an experiment where the UAV performed a 35-s flight incorporating linear and nonlinear motion states. The first 20 seconds involved uniform linear motion at a velocity of $2m/s$. Afterward, the UAV transitioned into a uniform circular motion mode, maintaining a linear velocity of $2m/s$ for the remaining 15 seconds. This flight trajectory was used for the ablation experiments.
6.3.1 Contributions of different attacks
To explore the contribution of velocity and positional attacks to overall attack effectiveness, we designed comparison experiments with three attack strategies: single position perturbation attack (SPA), single velocity perturbation attack (SVA), and combined concerted attack (CCA). The quantitative analysis of ADEs, presented in Table II, shows that both individual attacks are effective. Specifically, the ADEs for SPA and SVA are improved by 184% and 212%, respectively, compared to the baseline. When comparing the combined attack (CCA) to the single attacks, the ADE improves by 129% over SPA and 112% over SVA. This indicates that CCA not only significantly enhances attack effectiveness by leveraging the coupling effect between position and velocity, but also perturbs both longitude and latitude directions simultaneously. Furthermore, the stealth assessment results in Fig. 8 show that while the CCA approach induces brief oscillations in the detection statistics, these fluctuations remain well within the acceptable threshold limits, ensuring that the attack remains stealthy.
TABLE II: Ablation Study of Different Attack’s Contribution
Position N Position E
Baseline 1.41 0.66
SPA 2.72 0.89
SVA 2.45 2.2
CCA 2.79 2.41
Figure 8: $\chi^{2}$ comparation
6.3.2 Attack Combination
We apply the velocity attack during the linear motion state and the position attack during the nonlinear motion state and examine the impact of this combination on the attack’s stealthiness. The experimental results, shown in Fig. 9, reveal that $9s$ after the attack begins, the detector quickly identifies the existence of the attack. Furthermore, even when the UAV transitions to a nonlinear motion state, the detector continues to successfully detect the attack for $5s$, despite the change in attack mode. This phenomenon occurs because the drastic velocity changes in the linear motion state cause significant fluctuations in the residuals. In contrast, the nonlinear motion state allows SSD to effectively smooth out the impact of the velocity attack, ensuring the attack remains stealthy throughout the process. This further validates the rationale of the SSD framework.
Figure 9: Detection Statistic comparison for attack combination
6.4 Attack Stealthiness
To validate the stealthiness, we selected chi-square detector for attack detection. We also chose biased signal and multiplicative attacks for comparison and evaluated them across three mission trajectories. Based on prior research, we designed the following attack payloads: (1) GNSS positions are added with a uniform distribution U(0, 0.0005); (2) GNSS positions are scaled by a factor of 1.5. The experimental results in Fig. 10 demonstrate that SSD successfully limits detection statics to the threshold range and bypasses both detection methods with a carefully chosen set of attack parameters. Notably, compared to these works, we significantly reduced the loadings applied for the biased signal Attack. However, both detection methods were able to detect the attack effectively. It is clear that SSD exhibits strong stealthiness properties under chi-square detection methods and several motion states.
(a) $\chi^{2}$ detection of different attacks under Trajectory I
(b) $\chi^{2}detection$ of different attack under Trajectory II
(c) $\chi^{2}detection$ of different attack under Trajectory III
Figure 10: Attack stealthiness evaluation
We also use two UAV-specific detectors, NLC [30] and LTW [31], as baselines for comparison. These methods are highly effective at detecting GNSS attacks and they maintain both strong accuracy and response time, even when identifying targeted stealthy attacks. We follow the threshold in previous work [30]. Fig. 11 demonstrates that in the LTW test group, the detection statistic consistently remains below the threshold boundary, indicating the superior stealth characteristics of SSD attacks. Notably, NLC statistics demonstrate the following three characters (shown in Table III):
  • In Trajectory I (pure linear motion), the detection statistic approaches the threshold at $6s$ but never exceeds the threshold.
  • In Trajectory II (fully nonlinear motion), the detection statistic surpasses the threshold after $31s$ cumulative duration. The detection latency is largely increased compared to the previous work (about $0.3s$[30].
  • In Trajectory III (hybrid motion mode), no significant statistical fluctuations occur during $0-20s$ linear phase, with limited oscillations (peak statistic is 9.27) emerging post nonlinear component introduction at 20s.
TABLE III: Detection Performance Comparison for NLC
Detection Latency(s) Peak Statistic Success Rate
Trajectory I N/A 19.69 0%
Trajectory II 31s >$20$ 35.4%
Trajectory III N/A 9.27 0%
These findings demonstrate significant motion-dynamic sensitivity disparities in NLC detectors. However, when attackers combine motion-state transition strategies, positional attack can realize residual normalization during linear phases and reduce detection sensitivity velocity attack counteract statistical fluctuations from nonlinear components. These two methods make SSD show greater stealthiness in hybrid motion missions.
(a) Detection statistic in Trajectory I
(b) Detection statistic in Trajectory II
(c) Detection statistic in Trajectory III
Figure 11: Detection statistic comparison for NLC and LTW under SSD.
6.5 Attack Effectiveness
We start by determining the optimal combination of $Q_{k}$ and $R_{k}$ for each scenario through offline learning. With these optimal values, we proceed to validate the effectiveness of the SSD. For each combination of INS and trajectory, the UAV is tasked with following the designated path to complete a full mission, while the SSD is deployed to attack the flight. Table IV presents the changes in each metric before and after the attack. On average, ADE/FDE is increased by 425%/591%. The lateral (N)/longitude (E) deviation reaches / meters. We will analyze the factor that based on the experiment on three scenarios.
TABLE IV: Attack Effectiveness
Model Scenario Duration ADE FDE APDE
Normal/Attack Normal/Attack Normal/Attack
ECL EKF2 Trajectory I 20s (0.95)/(5.68) 597.89% (1.80)/(5.89) 327.22% (0.89)/(5.66) 635.96%
CD-EKF Trajectory I (1.62)/(4.64) 286.42% (1.49)/(7.54) 506.04% (1.81)/(4.92) 273.48%
ECL EKF2 Trajectory II 48s (1.31)/(4.09) 312.21 % (1.78)/(6.15) 345.51% (1.48)/(4.11) 277.70%
CD-EKF Trajectory II (1.78)/(8.53) 479.21 %
ECL EKF2 Trajectory III 71s (1.46)/(5.38) 366.44% (0.95)/(10.47) 1102.11% (1.86)/(6.49) 348.92%
CD-EKF Trajectory III (1.65)/(5.38) 326.06% (0.56)/(10.47) 1869.64
(a) Attack visualization in Trajectory I
(b) Attack visualization in Trajectory II
(c) route
Figure 12: Attack visualization in Trajectory III
6.5.1 Different Scenarios
In terms of scenarios, the SSD shows a greater increase in positioning error in the purely linear motion state (Fig. 12(a)) compared to the purely nonlinear state (Fig. 12(b)). This difference arises from the time-varying nature of the velocity vector in the nonlinear state, which leads to attenuation of the indirect positional interference caused by the velocity perturbation $G(t_{i},a^{d}_{i};\phi)$. However, experiments with Section 6.4 demonstrate that a velocity attack can still ensure the fulfillment of the stealthy precondition. When the mission involves mixed kinematic modes, the combination of velocity and positional attacks results in an additive interaction, as shown in Table IV. The baseline increase is 366.44% for ADE and 1102.11% for FDE, confirming that SSD enhances attack effectiveness synergistically. Next, we will quantitatively analyze the impact of this state on mission completion rates.
6.5.2 Impact on Mission
When the UAV’s trajectory deviation exceeds the tolerance range, the mission completion rate shows a clear decline. In three different scenarios, the APDE metrics increased by $635.96\%$, $277.70\%$, and $348.92\%$, respectively, indicating that SSDs significantly compromise mission reliability. Notably, in Scenario Trajectory III, where the mission duration is 71 seconds, the FDE and APDE deviations reached $10.47m$ and $6.49m$, respectively—the largest among the three scenarios. This demonstrates that SSD lowers the mission completion rate and impedes the UAV’s ability to return accurately after completing the mission (shown in Fig. 12(c)). This effect highlights its kinematic chain impact:since the mission-critical point is typically located at the motion state transition (e.g., hover → level flight, acceleration → deceleration). The attack-induced cumulative error in position estimation propagates into the next motion state, where it is further amplified.
7 Discussion
This study reveals the coupling mechanism between UAV motion dynamics and the cyber attacks effectiveness through systematic empirical analysis. Experimental data indicate that changes in motion state can significantly enhance the success rate of attacks. While the SSD approach demonstrates clear advantages, its engineering implementation faces two major challenges: Dependency on A Priori Knowledge. The effectiveness is highly contingent upon the real-time accuracy of the object detection and tracking system. The existing YOLOv5 architecture, for instance, exhibits exponential decay in the Intersection over Union (IoU) metric over time in dynamic target tracking scenarios. This results in a tracking failure probability exceeding 73% after 60 seconds of continuous locking. Energy-Concealment Trade-off Paradox. While the sustained attack mode can maintain a stealthiness threshold, it leads to a non-linear increase in energy consumption on the attacking end. This escalation doesn’t align with the requirements in real-world mission scenarios.
In addition, a key assumption of the SSD is that the $Q(k)$ and $R(k)$ are preset offline. This is a common practice in current engineering applications. From a theoretical standpoint, adaptive $Q(k)$ and $R(k)$ values can modify the system’s sensitivity to attacks, potentially enabling mitigation of such attacks. This insight provides a constructive direction for defending SSD. We observed that increasing $Q(k)$ and decreasing $R(k)$ could reduce the system’s sensitivity to attacks. However, this adjustment comes at the cost of a decrease in localization accuracy. While it is possible to improve positioning accuracy by tweaking these parameters, this also makes the system more vulnerable to attacks. Consequently, a dynamic strategy for adjusting $Q(k)$ and $R(k)$ is crucial. This can be achieved through optimization methods or reinforcement learning, which can fine-tune these parameters in real time, balancing between accuracy and security.
8 Conclusion
In this paper, we investigate cybersecurity threats of UAV route planning under different motion states. We assess the effectiveness of GNSS attacks under various motion states through theoretical and experimental analyses. Our findings reveal that INS is more vulnerable during maneuvering than in linear flight. Based on this insight, we introduce SSD, a novel state-based stealthy backdoor attack, which strategically combines GNSS velocity and position attacks to exploit this vulnerability. We conducted extensive experiments, and the results show that SSD demonstrates superior effectiveness and stealthiness compared with previous methods. We hope that this work will inspire INS designers and developers to prioritize code security and implement robust dynamic defenses.
Acknowledgments
This work is supported by is supported by the National Natural Science Foundation of China (No.62103330, No.62233014), andthe Innovation Foundation for Doctor Dissertation of Northwestern Polytechnical University(CX2023023).
*: Corresponding author